### v2

# SuppSupplementary References

Abstract
A principal curve serves as a powerful tool for uncovering underlying structures of data through 1-dimensional smooth and continuous representations. On the basis of optimal transport theories, this paper introduces a novel principal curve framework constrained by monotonicity with rigorous theoretical justifications. We establish statistical guarantees for our monotone curve estimate, including expected empirical and generalized mean squared errors, while proving the existence of such estimates. These statistical foundations justify adopting the popular early stopping procedure in machine learning to implement our numeric algorithm with neural networks. Comprehensive simulation studies reveal that the proposed monotone curve estimate outperforms competing methods in terms of accuracy when the data exhibits a monotonic structure. Moreover, through two real-world applications on future prices of copper, gold, and silver, and avocado prices and sales volume, we underline the robustness of our curve estimate against variable transformation, further confirming its effective applicability for noisy and complex data sets. We believe that this monotone curve-fitting framework offers significant potential for numerous applications where monotonic relationships are intrinsic or need to be imposed.
Keywords: Principal Curve, Statistical Error Analysis, Optimal Transport, Neural Network
1 Introduction
A principal curve is a 1-dimensional manifold penetrating the middle of data. In mathematics, a general curve is defined as a vector of functions where an index parameter draws the shape of each component function. For example, we can parametrize a unit circle as $(\cos s,\sin s)$, or a quadratic curve in $\mathbb{R}^{2}$ as a vector of $(s,s^{2})$, where $s\in{\mathbb{R}}$ denotes the index parameter. Similarly, a principal curve is a 1-dimensional manifold of multivariate random variables whose probability law relies on the index parameter. A representative example is a multivariate Gaussian distribution whose mean vector varies by the index parameter. In this regard, a principal curve captures the essential geometric shape of data, and it also provides curve-sense summary statistics to analyze joint relationships between variables rather than conditional relationships. The latter characteristic of a principal curve can be particularly beneficial when causal relationships between variables are ambiguous, which occurs a lot in numerous real-world problems.
Various fields harness such characteristics of a principal curve in data analysis. Banfield and Raftery (1992) clustered principal curves to outline distinctive floes in a satellite image as an image processing method. In the field of transportation, Einbeck and Dwyer (2011) analyzed the relationship between vehicles’ speed and traffic flow by curves since the two key variables are hardly seen to have a causal relationship with each other but rather a latent variable traffic density associates the target variables. Also, for pathway-level analysis on cancer progression, Drier et al. (2013) projected normal and tumor samples on a principal curve estimated from the principal components of the two contrasting samples and showed the distance along the curve between the samples reveals important findings on cancer data analysis. Chen et al. (2015) employed a principal curve technique to detect and compare cosmic filaments from densities of the galaxy and dark matter. For more examples of using principal curves for data analysis, we refer to Takei et al. (2021); Layton et al. (2022); Ghazanfar et al. (2024).
While the literature on principal curves becomes more mature, the concept of a principal curve constrained to exhibit a monotonic shape remains underexplored. Monotonic relationships between variables are prevalent in numerous real-world scenarios, addressing the practical importance of estimating a monotone principal curve for reliable analysis. For instance, demand and supply curves in economics illustrate inverse and direct relationships between price and quantity. Arfaoui and Ben Rejeb (2017) analyzed the interdependencies between oil, gold, US dollars, and the stock market through a simultaneous equation model, which justifies that each variable relates positively or negatively to other variables. As addressed in Patton and Timmermann (2010), a plethora of theories in financial modeling require monotonic situations such as discovering the relation between expected returns and market betas in the capital asset pricing model. In ethics-considered machine learning, such as for privacy or fairness, ethical improvement of a learned model tends to trade off its utility such as degree of accuracy in general (Xie et al., 2018; Sohn et al., 2024). These real-world problems highlight the potential utility of monotone principal curves as fundamental data analysis tools when monotonic phenomena between variables are expected.
Our ontributions are mainly threefold. First, we propose a novel principal curve framework for fitting a monotonic curve in Section 2. To begin with, we characterize the monotone curve in arbitrary finite-dimensional space using convex analysis and monotone operator theory, and then we propose an optimization problem to find the monotone curve. Secondly, Section 3 places rigorous theoretical analysis on the curve estimate, the solution of the established optimization. After verifying the existence of the solution, we disclose the convergence rate of the expected empirical and generalization errors. The generalization error in this work implies the estimation error for unseen data. Note that tackling such a generalization error is particularly important from machine learning and statistics viewpoints, since it can provide a theoretical foundation for model selection to avoid overfitting or find optimal hyperparameters. To our knowledge, this is the first work in the principal curve literature to derive the generalization error based on convex analysis. Finally, in Section 4, we introduce a numerical algorithm that finds the monotone curve through neural networks. In Section 5, simulation studies verify that our method achieves more accurate estimation than general curve-fitting methods when the underlying curve is monotone. Section 6 applies the proposed framework to model future prices for copper, gold, and silver, and to estimate a demand curve between avocado prices and sales volume. These real data applications advertise that our method enjoys more robust estimation in terms of variable transformation than the competing methods when monotone relationships are anticipated.
1.1 Literature review
To seek a better representation of a principal curve, versatile approaches have been steadily attempted and categorized mainly by two branches from our perspective. First, a principal curve is seen as a minimizer of a distance between the curve and data points with regularization. Secondly, a principal curve is approached by a ridge where gradients and eigenvectors of a density’s Hessian are orthogonal. The seminal work of Hastie and Stuetzle (1989) defined a curve on self-consistency as a local average of data points having the minimal mean squared distance to the data points indexed by the domain of the curve itself. This initial approach, however, guarantees the existence of a curve only on a specific data distribution and tends to estimate a biased curve around high-curvature areas. These limitations ignited subsequent studies. To alleviate the bias, Tibshirani (1992) approached a principal curve as weighted local averages of data points by formulating the curve-generating process as a probabilistic mixture modeling. Banfield and Raftery (1992) eased the bias by updating a curve estimate based on the smoothed residuals of the curve. To ensure the existence of a curve in a more general distribution, Kégl et al. (2000) considered a curve whose length is bounded by a fixed size on a bounded and convex domain. Gerber and Whitaker (2013) proposed a surrogate objective to replace the mean squared distance so that a principal curve lies in critical points that are locally minimal. In behind, Duchamp and Stuetzle (1996) showed the self-consistent curve can be critical but not extremal points for the mean squared distance, which contributed to justifying such challenges of estimating self-consistent curves. Recent efforts in this branch have focused on finding self-consistent curves in non-Euclidean spaces (Hauberg, 2015; Lee et al., 2020; Kang and Oh, 2024).
On the one hand, Ozertem and Erdogmus (2011) viewed a principal curve, for a given probability density function, to lie in an 1-dimensional intrinsic space in which the density’s gradient and eigenvectors of the Hessian are constrained to be orthogonal. The set of points in the constrained space is called a ridge or filament that captures the essential structures of the data (Hall et al., 1992; Eberly et al., 1994; Qiao and Polonik, 2016). In general, ridges refer to multidimensional constrained spaces where the projected gradient of the density vanishes in certain directions defined by the Hessian eigenvectors, and the curvature satisfies specific eigenvalue constraints. Note the filament stands for the 1-dimensional ridge. Genovese et al. (2014) mathematically justified that this subspace-constrained approach that finds a ridge can approximate the true lower-dimensional manifold with an additive noise term. Chen et al. (2015) suggested selecting a reasonable kernel’s bandwidth used to approximate the probability density by maximizing the extent of covering a population ridge. On the other hand, Qiao and Polonik (2016) viewed the 1-dimensional ridge as an integral curve from the differential equation involving the Hessian’s eigenvector corresponding to the second eigenvalue in the two-dimensional space. Besides, further efforts have been made to improve the algorithmic computation and theory of the subspace-constrained methodology (Qiao and Polonik, 2021; Zhang and Chen, 2023).
2 Formulation of monotone curve-fitting task
The construction of our monotone curve-fitting task starts with characterizing a monotone curve in a finite-dimensional space through convex analysis. Then a novel learning problem is introduced where the solution corresponds to the suggested monotone curve.
2.1 Monotone set and diagonal parametrization
Let $[n]:=\{1,2,...,n\}$ for $n\in{\mathbb{N}}=\{1,2,...\}$. Let ${\cal P}(\Omega)$ denote the set of probability measures (distributions) over a measure space $\Omega$, and $\langle{\bf a},{\bf b}\rangle$ denote the inner product of vectors ${\bf a},{\bf b}$. We say that a set $\Gamma\subset{\mathbb{R}}^{k}$ is monotone if for any ${\bf a}=(a_{i})_{i},{\bf b}=(b_{i})_{i}\in{\mathbb{R}}^{k}$, either $a_{i}\geq b_{i}$ or $a_{i}\leq b_{i}$ for all $i\in[k]$. Note that this is equivalent to $(b_{j}-a_{j})(b_{i}-a_{i})\geq 0$ for all $i,j\in[k]$. A monotone set $\Gamma$ is maximally monotone if it is not a proper subset of other monotone sets.
Now we characterize a monotone curve via the diagonal coordinate $s({\bf x}):=\sum_{i=1}^{k}x_{i},\forall{\bf x}=(x_{1},...,x_{k})\in{\mathbb{R}}^{k}$. Let $S(\Gamma):=\{s({\bf x})\,|\,{\bf x}\in\Gamma\}$ and $s_{\Gamma}:\Gamma\to S(\Gamma)$, the restriction of $s$ on $\Gamma$.
Definition 2.1
For a monotone set $\Gamma\subseteq{\mathbb{R}}^{k}$, we call $\gamma=(\gamma_{1},...,\gamma_{k}):=s_{\Gamma}^{-1}:S(\Gamma)\to\Gamma$ the associated monotone curve, which parametrizes $\Gamma$ via the diagonal coordinate $s$.
Note that $s_{\Gamma}$ is bijective and each component $\gamma_{i}$ is a nondecreasing function of $s$ if $\Gamma$ is monotone, and the maximal monotonicity of $\Gamma$ is equivalent to the condition $S(\Gamma)={\mathbb{R}}$. Our definition of a monotone curve is based on an underlying monotone set. In the following, we outline the characterization of a monotone set using convex functions that satisfy a duality relationship, and show that the associated curve $\gamma$ can be represented by them.
2.2 Exposure of a monotone set through convex functions
We first introduce basic mathematical tools for convex analysis. Let ${\cal A}({\cal H})$ denote the set of proper, lower-semicontinuous and convex functions (valued in {${\mathbb{R}}\cup\{+\infty\}$) on a Hilbert space ${\cal H}$. In this general setting, a set $\Gamma\subset{\cal H}\times{\cal H}$ is called monotone if $\langle b_{2}-a_{2},b_{1}-a_{1}\rangle\geq 0$ for any ${\bf a}=(a_{1},a_{2}),{\bf b}=(b_{1},b_{2})\in\Gamma$. For a function $f:{\cal H}\to{\mathbb{R}}\cup\{+\infty\}$, its convex conjugate is $f^{*}({\bf y}):=\sup_{{\bf x}\in{\cal H}}\,[\langle{\bf x},{\bf y}\rangle-f({\bf x})]$. Then the Fenchel-Young inequality $f({\bf x})+f^{*}({\bf y})\geq\langle{\bf x},{\bf y}\rangle$ holds for all ${\bf x},{\bf y}\in{\cal H}$. We say that $f$ and $g$ are mutually conjugate if $f=g^{*}$ and $g=f^{*}$. The Fenchel–Moreau theorem states that $f^{**}:=(f^{*})^{*}=f$ for any $f\in{\cal A}({\cal H})$, which implies $f$ and $f^{*}$ are mutually conjugate. For $f\in{\cal A}({\cal H})$ and ${\bf x}\in{\cal H}$, the subdifferential of $f$ at ${\bf x}$ is defined as the following convex set ${\partial}f({\bf x})=\{{\bf v}\in{\cal H}\mid f({\bf y})-f({\bf x})\geq\langle{\bf v},{\bf y}-{\bf x}\rangle\,\forall{\bf y}\in{\cal H}\}$, and the set ${\partial}f=\bigcup_{{\bf x}\in{\cal H}}\{({\bf x},{\bf y})\in{\cal H}\times{\cal H}\mid{\bf y}\in{\partial}f({\bf x})\}$ is then called the subdifferential of $f$.
The following proposition by Rockafellar and Minty (see Bauschke and Combettes (2019)) states that a monotone set is contained in a set where the Fenchel-Young inequality achieves equality. Let $S(\Gamma)=\{{\bf x}+{\bf y}\mid({\bf x},{\bf y})\in\Gamma\}$ and $\Gamma^{-1}:=\{({\bf y},{\bf x})\mid({\bf x},{\bf y})\in\Gamma\}$ for $\Gamma\subset{\cal H}^{2}$.
Proposition 2.1
Let $f,g\in{\cal A}({\cal H})$ satisfy $f(x)+g(y)\geq\langle x,y\rangle$ for all $x,y\in{\cal H}$. Then the contact set $\Gamma_{f,g}:=\{(x,y)\in{\cal H}^{2}\ |\ f(x)+g(y)=\langle x,y\rangle\}$ is monotone. Moreover, $f,g$ are mutually conjugate if and only if $\Gamma_{f,g}$ is maximally monotone if and only if $S(\Gamma_{f,g})={\cal H}$, in which case $\Gamma_{f,f^{*}}={\partial}f=({\partial}f^{*})^{-1}$. Furthermore, any monotone set $\Gamma\subset{\cal H}^{2}$ is contained in a maximally monotone set $\Gamma_{f,f^{*}}$ for some $f\in{\cal A}({\cal H})$.
The equality ${\partial}f=({\partial}f^{*})^{-1}$ states that for $f\in{\cal A}({\cal H})$, ${\partial}f$ and ${\partial}f^{*}$ are the (generalized) inverses of each other.
Understanding how to obtain such $f$ and $f^{*}$ that expose a given monotone set $\Gamma$ as $\Gamma\subset\Gamma_{f,f^{*}}$ is particularly simple and instructive when ${\cal H}={\mathbb{R}}$. To simplify discussion, we suppose that $\Gamma\subset{\mathbb{R}}^{2}$ is monotone and $\Gamma$ is maximally monotone without loss of generality.
Remark 2.1 (Construction of $f,f^{*}$ given $\Gamma$)
[…]
2.3 Loss function to extract a monotone set
Such procedures for characterizing a monotone set through mutually conjugate convex functions lay the groundwork for designing a novel statistical learning framework to identify a monotone principal curve. It is helpful to first discuss this problem in the two-dimensional space ${\mathbb{R}}^{2}$. Let $\mu\in{\cal P}({\mathbb{R}}^{2})$ denote the data distribution of ${\bf X}=(X_{1},X_{2})$, which exhibits an approximately monotone structure. To formulate a learning framework, this work proposes the following “loss function” to identify a monotone set:
$H({\bf x};f_{1},f_{2})=f_{1}(x_{1})+f_{2}(x_{2})-x_{1}x_{2},\quad{\bf x}=(x_{1},x_{2})\in{\mathbb{R}}^{2}$
where $f_{1}$ and $f_{2}$ are in duality defined as below.
Definition 2.2
We say that a pair of convex functions $f_{1},f_{2}\in{\cal A}({\mathbb{R}})$ is in duality position, or simply in duality, if $H({\bf x};f_{1},f_{2})\geq 0$, i.e., $f_{1}(x_{1})+f_{2}(x_{2})\geq x_{1}x_{2}$ for all $x_{1},x_{2}\in{\mathbb{R}}$.
This loss function will be generally larger as ${\bf x}$ is farther from the monotone set $\Gamma_{f_{1},f_{2}}$ on which $H$ attains its minimum 00. Figure 1 draws a contour of $H({\bf x};f,f^{*})$ for a power function on the grid of $x_{1},x_{2}$.
[…]
This inspires us to use ${\mathbb{E}}_{{\bf X}\sim\mu}[H({\bf X};f_{1},f_{2})]$ as an objective and minimize it over the convex pair $f_{1},f_{2}$ in duality, since Proposition 2.1 shows that the zero set $\Gamma_{f_{1},f_{2}}$ will then represent a monotone curve in $\mathbb{R}^{2}$. Later, we will present a generalized formulation for higher dimensions $\mathbb{R}^{k}$ in (2) and (3).
We now explain how the diagnoal parametrization (monotone curve) $\gamma=s_{\Gamma}^{-1}:S(\Gamma)\to\Gamma$ can be represented by a convex pair $(f_{1},f_{2})$. If $(f_{1},f_{2})$ is in duality and ${\bf x}=(x_{1},x_{2})$, the first order condition yields the following implication
$H({\bf x};f_{1},f_{2})=0\implies x_{2}\in{\partial}f_{1}(x_{1})\,\text{ and }\,x_{1}\in{\partial}f_{2}(x_{2}).$ (1)
With $q(x):=|x|^{2}/2$ and $g_{i}:=f_{i}+q$,  $i=1,2$, the following equivalences $x_{2}\in{\partial}f_{1}(x_{1})\iff x_{1}+x_{2}\in{\partial}g_{1}(x_{1})$ and $x_{1}\in{\partial}f_{2}(x_{2})\iff x_{1}+x_{2}\in{\partial}g_{2}(x_{2})$ are straightforward. Hence, the implication (1) can be restated as ${\bf x}\in\Gamma_{f_{1},f_{2}}\implies x_{i}=({\partial}g_{i})^{-1}(s({\bf x})),\ i=1,2$. In fact, the reverse implication also holds; see (Bartz et al., 2021, Theorem 4.1) and its proof.
This shows that the diagonal parametrization $\gamma=(\gamma_{1},\gamma_{2})$ of the monotone set $\Gamma_{f_{1},f_{2}}$ can be expressed as $\gamma_{i}=({\partial}g_{i})^{-1}$. Moreover, $({\partial}g_{i})^{-1}=\nabla g_{i}^{*}$, where (whether $g_{i}\in{\cal A}({\mathbb{R}})$ is differentiable or not) $g_{i}^{*}=f_{i}^{*}\square q:=\inf_{y\in{\mathbb{R}}}[f_{i}^{*}(y)+q(x-y)]$, called the Moreau envelope of $f_{i}^{*}$, is a differentiable convex function on ${\mathbb{R}}$ with $1$-Lipschitz nondecreasing derivative $\nabla g_{i}^{*}$ (Bauschke and Combettes (2019)). Hence the monotone curve $\gamma(s):=\big{(}({\partial}g_{1})^{-1}(s),({\partial}g_{2})^{-1}(s)\big{)}$ is defined for all $s\in{\mathbb{R}}$, with $\big{(}({\partial}g_{1})^{-1}(s),({\partial}g_{2})^{-1}(s)\big{)}\in\Gamma_{f_{1},f_{2}}$ if $s\in S(\Gamma_{f_{1},f_{2}})$. In particular, if $f_{1}=f=f_{2}^{*}$, then since $S(\Gamma_{f,f^{*}})={\mathbb{R}}$, $\gamma$ yields a bijection between ${\mathbb{R}}$ and $\Gamma_{f,f^{*}}$.
2.4 Exposure of a monotone curve in higher dimensions
This section extends the aforementioned discussion to any data dimension $k$. For $f_{i}\in{\cal A}({\mathbb{R}})$, $i\in[k]$, we write ${\bf f}=(f_{1},...,f_{k})$ and $f^{\oplus}({\bf x}):=\sum_{i=1}^{k}f_{i}(x_{i})$ for ${\bf x}=(x_{1},...,x_{k})\in{\mathbb{R}}^{k}$. Define a cost function $c({\bf x}):=\sum_{1\leq i<j\leq k}x_{i}x_{j}$ and $H({\bf x};{\bf f}):=f^{\oplus}({\bf x})-c({\bf x})=\sum_{i=1}^{k}f_{i}(x_{i})-\sum_{1\leq i<j\leq k}x_{i}x_{j}$. Let us say that ${\bf f}$ is in duality position, or simply in duality, if the nonnegativity $H({\bf x};{\bf f})\geq 0$ holds for all ${\bf x}\in{\mathbb{R}}^{k}$. Also we say that ${\bf f}=(f_{1},\ldots,f_{k})$ is a $c$-conjugate tuple if for each $i\in[k]$ and $x_{i}\in{\mathbb{R}}$, $f_{i}(x_{i})=(\bigoplus_{j\neq i}f_{j})^{c}(x_{i}):=\sup_{j\neq i,\,x_{j}\in{\mathbb{R}}}\ [c(x_{1},\dots,x_{i},\dots,x_{k})-\sum_{j\neq i}f_{j}(x_{j})]$. Any $c$-conjugate tuple is in duality position. Also, $(\bigoplus_{j\neq i}f_{j})^{c}$ is convex and lower-semicontinuous for any proper (not necessarily convex) functions $(f_{j})_{j}$, thus $(\bigoplus_{j\neq i}f_{j})^{c}\in{\cal A}({\mathbb{R}})$ if proper.
Remark 2.2
Let $q(x)=x^{2}/2$, $g_{i}=f_{i}+q$. We then have $H({\bf x};{\bf f})=\widetilde{H}({\bf x};{\bf g})$ with $\widetilde{H}({\bf x};{\bf g}):=g^{\oplus}({\bf x})-\tilde{c}({\bf x})$ and $\tilde{c}({\bf x}):=s({\bf x})^{2}/2$. As a result, ${\bf f}$ is $c$-conjugate if and only if ${\bf g}$ is $\tilde{c}$-conjugate.
The following is an extension of Proposition 2.1 for the $c$-conjugate functions on ${\mathbb{R}}$.
Proposition 2.2
Assume that ${\bf f}=(f_{i})_{i=1}^{k}\in{\cal A}({\mathbb{R}})^{k}$ is in duality position. Then the zero set $\Gamma_{{\bf f}}:=\{{\bf x}\in{\mathbb{R}}^{k}\,|\,H({\bf x};{\bf f})=0\}$ is monotone. Moreover, ${\bf f}$ is a $c$-conjugate tuple if and only if $\Gamma_{{\bf f}}$ is maximally monotone if and only if $S(\Gamma_{{\bf f}})={\mathbb{R}}$. Furthermore, any monotone set $\Gamma\subset{\mathbb{R}}^{k}$ is contained in a maximally monotone set $\Gamma_{{\bf f}}$ for some $c$-conugate tuple ${\bf f}\in{\cal A}({\mathbb{R}})^{k}$.
The monotonicity of $\Gamma_{{\bf f}}$ is shown in Lemma 8.1 in Supplementary 8. The proposition shows that any monotone curve in ${\mathbb{R}}^{k}$ can be exposed as a subset of the monotone set $\Gamma_{{\bf f}}$ induced by convex functions ${\bf f}$ in duality, and $\Gamma_{{\bf f}}$ is maximally monotone if and only if ${\bf f}$ is $c$-conjugate. This motivates us to use $H({\bf x};{\bf f})$ as the objective function. We note that Remark 2.1 can also be extended for ${\mathbb{R}}^{k}$. This is outlined in Supplementary 7.
To characterize a monotone curve $\gamma$ in ${\mathbb{R}}^{k}$, we turn to the parametrization of monotone sets in ${\mathbb{R}}^{k}$ using the diagonal coordinate. Given $f_{i}\in{\cal A}({\mathbb{R}})$, $i\in[k]$, with $H(\,\cdot\,;{\bf f})\geq 0$ and ${\bf x}=(x_{1},...,x_{k})\in{\mathbb{R}}^{k}$, the equivalence $H({\bf x};{\bf f})=0\iff s({\bf x})\in{\partial}g_{i}(x_{i})\ \forall i\in[k]$ holds true by an application of Theorem 4.1 of Bartz et al. (2021). This yields a parametrization of the monotone set $\Gamma_{{\bf f}}$ via the diagonal coordinate, i.e., ${\bf x}\in\Gamma_{{\bf f}}\iff x_{i}=({\partial}g_{i})^{-1}(s({\bf x})),\ \forall i\in[k]$, where $({\partial}g_{i})^{-1}=\nabla g_{i}^{*}=\nabla(f_{i}^{*}\square q)$ is a $1$-Lipschitz nondecreasing function on ${\mathbb{R}}$. We can thus write $\gamma(s)=\gamma_{{\bf f}}(s):=\big{(}({\partial}g_{i})^{-1}(s)\big{)}_{i\in[k]}$. If ${\bf f}$ is $c$-conjugate, then $\gamma:{\mathbb{R}}\to\Gamma_{{\bf f}}$ is a bijection.
2.5 Monotone curve-fitting task
The diagonal parametrization motivates another natural form for the objective function $\|{\bf x}-\gamma_{{\bf f}}(s({\bf x}))\|^{2}$. Intuitively, $\gamma_{{\bf f}}$ minimizing the squared error can be seen as a 1-dimensional latent manifold that condenses the original information of the data points. This extra term, therefore, helps $\gamma_{{\bf f}}$ represent the essential geometric shape of the data points. In manifold learning, this term is commonly referred to as a reconstruction error, representing the extent to which a model can accurately reproduce the original data. Consequently, we propose the following optimization problem for the monotone curve-fitting task (for a chosen parameter $\lambda\geq 0$ and a domain ${\cal D}\subset{\cal A}({\mathbb{R}})$):
minimize $\text{minimize }\,{\mathbb{E}}_{{\bf X}\sim\mu}\big{[}H({\bf X};{\bf f})+\lambda\|{\bf X}-\gamma_{{\bf f}}(s({\bf X}))\|^{2}\big{]}\text{ over }{\bf f}=(f_{1},...,f_{k})\in{\cal D}^{k}\text{ in duality.}$. (2)
As before, if we restrict ${\bf f}$ to be $c$-conjugated, then the problem is expressed as
minimize $\text{minimize }\,{\mathbb{E}}_{{\bf X}\sim\mu}\big{[}H({\bf X};{\bf f})+\lambda\|{\bf X}-\gamma_{{\bf f}}(s({\bf X}))\|^{2}\big{]}\text{ over $c$-conjugate tuples}\ {\bf f}\in{\cal D}^{k}.$ (3)
As a result, the solution curve $\gamma_{{\bf f}}$ becomes restricted to be monotone in $\Gamma_{{\bf f}}$ while having a minimal distance to the data.
This penalty-based optimization categorizes our monotone curve-fitting task to the first branch in the curve literature discussed in Section 1.1. More specifically, the self-consistency curve $\gamma_{\text{SC}}$ of Hastie and Stuetzle (1989) can also be seen to minimize the reconstruction error ${\mathbb{E}}[\|{\bf X}-\gamma_{\text{SC}}(s)\|^{2}]$ in the sense that their approach is to find the projection index that has smaller reconstruction error $s_{\gamma_{\text{SC}}}({\bf X}):=\sup\{s:\|{\bf X}-\gamma_{\text{SC}}(s)\|=\inf_{s^{\prime}}\|{\bf X}-\gamma_{\text{SC}}(s^{\prime})\|\}$ where $\gamma_{\text{SC}}(s)={\mathbb{E}}[\,{\bf X}\,|\,s_{\gamma_{\text{SC}}}({\bf X})=s\,]$. In our case, we specify the diagonal coordinate, which is not a variable of optimization, but still maintains the data-driven nature by summing all variables. Similarly, other methodologies in this branch formulate constrained optimization to define and find principal curves (Tibshirani, 1992; Kégl et al., 2000; Hauberg, 2015).
In conclusion, we propose the following monotone curve-fitting task.
Task 1
Given data $\widehat{\mu}\in{\cal P}({\mathbb{R}}^{k})$, solve (2) or (3) to find a minimizer of convex functions $\widehat{{\bf f}}=(\widehat{f}_{1},\dots,\widehat{f}_{k})$. The parametrized monotone curve $\widehat{\gamma}(s):=\big{(}({\partial}\widehat{g}_{i})^{-1}(s)\big{)}_{i\in[k]}$ is offered as a solution for the monotone curve-fitting task, where $\widehat{g}_{i}=\widehat{f}_{i}+q$.
2.5.1 Extension of Task 1 via orthogonal transformation
What can we do if, for instance, the observed data in ${\mathbb{R}}^{2}$ does not exhibit rough monotonicity but instead aligns with an anti-monotone structure, such as along the anti-diagonal? In cases where the data appears to be roughly aligned along an axis that is not parallel to the diagonal, one could consider applying Task 1 after performing an appropriate rotation of the data. We denote by ${\cal O}(k)=\{U\in{\mathbb{R}}^{k\times k}\,|\,U^{\mathsf{T}}U=UU^{\mathsf{T}}=I\}$ the set of all orthogonal matrices, where $U^{\mathsf{T}}$ is the transpose of $U$ and $I$ is the identity matrix. Then for ${\bf f}=(f_{1},\dots,f_{k})\in{\cal D}^{k}$ with a chosen domain ${\cal D}\subset{\cal A}({\mathbb{R}})$, the problems (2) and (3) can be generalized as
minimize $\text{minimize }{\mathbb{E}}_{{\bf X}\sim\mu}\big{[}H(U{\bf X};{\bf f})+\lambda\|U{\bf X}-\gamma_{{\bf f}}(s(U{\bf X}))\|^{2}\big{]}\text{ over $U\in{\cal O}(k)$ and ${\bf f}$ in duality},$ (4)
minimize $\text{minimize }{\mathbb{E}}_{{\bf X}\sim\mu}\big{[}H(U{\bf X};{\bf f})+\lambda\|U{\bf X}-\gamma_{{\bf f}}(s(U{\bf X}))\|^{2}\big{]}\text{ over $U\in{\cal O}(k)$ and $c$-conjugate }{\bf f},$ (5)
respectively. Consequently, we propose the following modified task.
Task 2
Given data $\widehat{\mu}\in{\cal P}({\mathbb{R}}^{k})$, solve (4) or (5) to find a minimizer of an orthogonal transformation $\widehat{U}$ and convex functions ${\widehat{\bf f}}=(\widehat{f}_{1},\dots,\widehat{f}_{k})$. The parametrized monotone curve $s\mapsto\widehat{U}^{\mathsf{T}}\widehat{\gamma}(s)$, where $\widehat{\gamma}(s)=\big{(}({\partial}\widehat{g}_{i})^{-1}(s)\big{)}_{i\in[k]}$ and $\widehat{g}_{i}=\widehat{f}_{i}+q$, is offered as a solution.
To motivate this task, we illustrate Task 2 on a toy example in Figure 2. A toy data set follows a bivariate normal variable ${\bf X}(s)=(X_{1}(s),X_{2}(s))^{\mathsf{T}}$ where ${\mathbb{E}}[X_{1}(s)]=-s^{2}$, ${\mathbb{E}}[X_{2}(s)]=\log(s+1)$, $\operatorname*{{Var}}(X_{1}(s))=\operatorname*{{Var}}(X_{2}(s))=0.1$, and $\operatorname*{{Cov}}(X_{1}(s),X_{2}(s))=0.09$ for $s\sim{\rm Unif}[0,3]$. Note that the ground truth curve of ${\bf X}$ is expressed by $(-s^{2},\log(s+1))^{\mathsf{T}}$. The original data points are positioned antidiagonally (left-top figure), which causes the transformation $U$ to align the transformed data points diagonally. On top of that, the contour of $H$ exposes the monotone set. Once the task finds the monotone curve, the estimated curve in the space of $U{\bf X}$ is rotated back to estimate the curve in the space of ${\bf X}$. Depending on a goal of an application, this final step may not be necessary if one wants to find a monotone curve on the space of $U{\bf X}$ or find $U$ to make variables monotonically related. A numeric algorithm to implement Task 2 appears in Section 4 later, where we use the first principal components of ${\bf X}$ to determine the initial value of $U$ and optimize it with other variables for sufficient flexibility.
Figure 2: Task 2 for the toy example: $(p_{1},p_{2})$ is the first principal component, $(1,1)$ indicates the diagonal direction, and $\gamma_{{\bf X}}$ is the true curve associated to ${\bf X}$.
3 Theoretical analysis
In this section, we first show the existence of our curve estimate found from the statistical learning problem and then investigate statistical error bounds for the estimate. The proof and the explicit forms of technical constants appear in Supplementary 8. To highlight the theoretical contributions of our work, we first review key literature in the field. We use the asymptotic notation $a_{n}=O(b_{n})$ (and $a_{n}=O_{p}(b_{n})$) to imply that $a_{n}/b_{n}$ is bounded above by a constant (and with high probability), as $n$ gets larger for the sequences $a_{n}$, $b_{n}$.
In the first branch of principal curve literature (see Section 1.1), the work of Kégl et al. (2000), assuming the curve has a fixed length on a bounded domain, showed that the rate of convergence is at $O(n^{-1/3})$, but the convergence is in terms of a loss function between data and a curve. That is, the loss function evaluating their empirical minimizer converges to the one of a true curve. More rigorous analysis of the statistical error rate stemmed from the second branch that defines the ridge of the target density as the principal curve. Genovese et al. (2014) provided statistical foundations for Ozertem and Erdogmus (2011). They found a ridge estimator from kernel density estimation converges to the population ridge at $O_{p}((\log n/n)^{2/(d+8)})$ in the Hausdorff distance, where $d$ is the dimension of data. Moreover, they showed that if data is concentrated on a manifold with Gaussian noise having standard deviation $\sigma$ in the ambient space, and the ridge estimator aims to extract the manifold, then the rate has an additive error term $O(\sigma^{2}\log(1/\sigma))$, relying on the size of the noise, that does not decay by the sample size.
3.1 Existence
To begin with, we show the existence of a solution to the problems (2) and (3) for any distribution of ${\bf X}\sim\mu$ where it has the finite second moment.
Theorem 3.1
Let ${\cal D}={\cal C}_{{\bf m}}$ in (2) and (3) where ${\bf m}=(m_{0},m_{1},m_{2},m_{3})$ with $m_{i}\geq 0$, and
[…]
$|f^{\prime\prime}(x)-f^{\prime\prime}(x^{\prime})|\leq m_{3}|x-x^{\prime}|\,\text{ for all }\,x,x^{\prime}\in{\mathbb{R}}\}.$
If ${\mathbb{E}}_{{\bf X}\sim\mu}\|{\bf X}\|^{2}<\infty$ and $m_{2}\geq k-1$, then solutions to (2) and (3) exist, respectively.
We remark that existing studies of a principal curve particularly for the first branch showed the existence of their principal curves under relatively stronger conditions. The principal curve with self-consistency condition in Hastie and Stuetzle (1989) guarantees the existence for some specific distributions such as spherically symmetric and ellipsoidal distributions. In Kégl et al. (2000), they showed the existence of a principal curve with a finite second moment, but they further require the curve to have a fixed length on a bounded domain. In contrast, our curve estimation enjoys the existence in more general distributions.
3.2 Statistical error analysis
In this section, we present our primary theoretical results on the statistical error bound of our estimate. We first outline the population model that generates the data. Consider a dimension $k\geq 2$, a maximally monotone curve $\Gamma\subset\mathbb{R}^{k}$, and a probability distribution $\rho\in\mathcal{P}(\mathbb{R}^{k})$ such that $\rho(\Gamma)=1$. Let $\varepsilon\in\mathcal{P}(\mathbb{R}^{k})$ denote the noise distribution with zero mean. Define ${\bf U}$ and ${\bf R}$ as independent random variables in $\mathbb{R}^{k}$, where ${\bf U}\sim\rho$ and ${\bf R}\sim\varepsilon$, thus ${\bf U}\in\Gamma$ and ${\mathbb{E}}[{\bf R}]=\textbf{0}$. The population model is then described by ${\bf X}={\bf U}+{\bf R}$ with ${\bf X}\sim\mu$, representing the population distribution. Let $\gamma(s)=\big{(}(\partial g_{i})^{-1}(s)\big{)}_{i\in[k]}$ parametrize $\Gamma$, where ${\bf f}=(f_{1},\dots,f_{k})$ is $c$-conjugate and exposes $\Gamma$, defined as in (9) in Supplementary 7.
Denote ${\bf X}=(X_{1},\dots,X_{k})$ and define $\widehat{\bf X}:=({\bf X}_{m})_{m=1}^{n}=(X_{1,m},\dots,X_{k,m})$, $m=1,\dots,n$, as independent and identically distributed (i.i.d.) observations sampled from $\mu$. The empirical data distribution is denoted by $\widehat{\mu}=\sum_{m=1}^{n}\delta_{{\bf X}_{m}}/n$, where $\delta_{x}$ denotes the Dirac mass at $x$. Let $\widehat{{\bf f}}\in{\cal D}^{k}$ be a $c$-conjugate solution to (3) with $\mu$ replaced by $\widehat{\mu}$, and let $\widehat{\gamma}(s)=\big{(}(\partial\widehat{g}_{i})^{-1}(s)\big{)}_{i\in[k]}$ denote the corresponding parametrized monotone curve. Note that although $\widehat{\mu}$ and $\widehat{\gamma}$ depend on the number of samples $n$, we omit it from the notations for simplicity.
We first establish an upper bound for the expected empirical MSE, defined as $E_{n}^{\textup{{emp}}}={\mathbb{E}}\left[n^{-1}\sum_{m=1}^{n}\|\gamma(s({\bf X}_{m}))-\widehat{\gamma}(s({\bf X}_{m}))\|^{2}\right]$, as stated in the following theorem. This result indicates that our curve estimate becomes more accurate as the noise in the data decreases.
Theorem 3.2
[…]